In a two-digit number, the digit in the tens place is three times the digit in the ones place. The sum of the digits is four more than the difference of the digits. What is the number?
Answer by stanbon(75887)

In a two-digit number, the digit in the tens place is three times the digit in the ones place. The sum of the digits is four more than the difference of the digits. What is the number?
-----------
Let the number be 10t+u
-----
Equations:
t = 3u
t+u = t-u + 4
------
Substitute for "t" and solve for "u"::
4u = 2u+4
2u = 4
u = 2
---
t = 3u = 6
---
Ans: 10t+u = 62
------------------
Cheers,
Stan H.
---------------